OpenStudy (anonymous):

FAN & MEDAL!!! HELP PLEASE!!! what is the equation of a vertical line passing through the point (-4,7) y=7 y=-4 x=7 x=-4

OpenStudy (anonymous):

@ShadowLegendX @sleepyjess @Coolsector

OpenStudy (sleepyjess):

The equation of a vertical line is x=x coordinate of the point

OpenStudy (anonymous):

so its x=-4? right so D?

OpenStudy (sleepyjess):

Yep

OpenStudy (anonymous):

thank you so much can you please help me with more so ik im doing them right?(:

OpenStudy (anonymous):

Given that f(x) = 4x − 3 and g(x) = quantity of 2x minus 1, divided by 3, solve for g(f(2)). 1 3 9 15

OpenStudy (anonymous):

@sleepyjess you there?

OpenStudy (anonymous):

can you help me please @AaronAndyson

OpenStudy (aaronandyson):

to find g(f(2)) first find f(2) f(x) is "short" for 4x-3 f(2) means use the formula 4x-3 but with x replaced with a 2 (and 4x means 4*x) can you find f(2) ?

OpenStudy (aaronandyson):

f(2) means 4*2 - 3 what do you get when you simplify that ?

OpenStudy (anonymous):

4*2=8 8-3=5 so the answer I got when I simplify it is 5

OpenStudy (aaronandyson):

correct :)

OpenStudy (anonymous):

im lost :( now

OpenStudy (aaronandyson):

sorry wait ok, now you know f(2) is 5 that means g( f(2) ) is the same as g(5) now we do the same thing, but use the formula for g g(x)=2x-1/3 can you find g(5) ?

OpenStudy (aaronandyson):

remember g(5) means replace x with 5: g(5)=2 times 5-1/3 what is 2 times 5 ? then subtract 1 that will be the "top" all divided by 3

OpenStudy (anonymous):

2*5=10 10-1=9 9/3=3 so the answer is 3 I got (:

OpenStudy (aaronandyson):

that's correct :)

OpenStudy (anonymous):

thank you so much may you please check/help me with more!!

OpenStudy (aaronandyson):

do so in a new post if i am unable to help other might help :)